compare the graph of each function with the graph of f(x)= x^2
1.g(x)=x^2-2
2.g(x)=2\3x^2
3.g(x)=5x^2+3
4.g(x)=-x^2+4
can u please help. Thank You

Answer by Alan3354(69443)

compare the graph of each function with the graph of f(x)= x^2
1.g(x)=x^2-2
2.g(x)=2\3x^2
3.g(x)=5x^2+3
4.g(x)=-x^2+4
--------------
dl the FREE graph software at
http://www.padowan.dk
------
Use Insert, then the function